Please help me solve the linear system by giving me the cordinates
X-5y =-3
3x - 5y =11

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

X-5y =-3
3x - 5y =11
----------------------
Subtract and solve for "x"::
2x = -8
x = -4
---------
Solve for "y"::
x - 5y = -3
-4 - 5y = -3
5y = -4+3
y = -1/5
--------------
Cheers,
Stan H.
----------------

Answer by Alan3354(69443)

X - 5y =-3
3x- 5y =11
================= Subtract